MsBrains (ms-brains):

What is the solution to the system of equations use the linear combination method 4x+y=5 3x+y =3 A.(-3,7) B.(2,-3) C.(0,-0) D.(0,5)

MsBrains (ms-brains):

@AloneS Can you help?

OpenStudy (anonymous):

I can

MsBrains (ms-brains):

really! Awesome!! Thank you

OpenStudy (anonymous):

We can always plug in the answers as they are in (x, y) format.

OpenStudy (anonymous):

Or I like to do this.

OpenStudy (anonymous):

NVM

MsBrains (ms-brains):

o.O

OpenStudy (anonymous):

Hold on, it does not work like that, it's not simple arithmetics....

OpenStudy (anonymous):

It's B, but I don't how to explain rather than plug in values...

OpenStudy (anonymous):

Oh I get it.

OpenStudy (anonymous):

4x+y=5 -3x-y =-3 Thus by subtracting literally from equations we get x=2

MsBrains (ms-brains):

Are you sure it's B?

OpenStudy (anonymous):

Such as 4x-3x+y-y=5-2

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

Literally subtract the two equations.

MsBrains (ms-brains):

Okay. I'm trusting you on this, if it's wrong, I'm going to kill you. .-. xD JK

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

If I am wrong, then math is a lie......

MsBrains (ms-brains):

I have 3 more

OpenStudy (anonymous):

It's 0 pretty sure.

OpenStudy (anonymous):

As you multiply 1.5 to the bottom equation and then subtract you get 0=6.5 which is not true.

OpenStudy (anonymous):

Reward me and then I will continue, xD

MsBrains (ms-brains):

Pfft, :P I did!

OpenStudy (anonymous):

Anymore, I solved two, I got two left.

MsBrains (ms-brains):

Yes, I have some gimme 1 sec

OpenStudy (anonymous):

That is 0 also..... 0=5?, that's a joke.

OpenStudy (anonymous):

BTW i am talking to myself.

MsBrains (ms-brains):

You are not! I'm just doing two things at once ^-^ sorry, I'll pay attention more.

OpenStudy (anonymous):

e=0....

MsBrains (ms-brains):

So there is no solution?

OpenStudy (anonymous):

Thus it is D or the last answer choice.

OpenStudy (anonymous):

I believe so, hold on, let me check.

MsBrains (ms-brains):

I have 1 last one. Can you help with it?

OpenStudy (anonymous):

Sure and I was right about it being D.

OpenStudy (anonymous):

Pretty sure, it's the intersection which I believe is (3, 0)

MsBrains (ms-brains):

Awesome. Thanks buddy!

OpenStudy (anonymous):

Your welcome.